# Pixel Reasoner: Incentivizing Pixel-Space Reasoning with Curiosity-Driven Reinforcement Learning

### v2

@@ Abstract
00footnotetext: Corresponding to: jasper.whz@outlook.com, dlwlrma314516@gmail.com, wenhuchen@uwaterloo.ca
@@ 1 Introduction
Recent advancements have demonstrated remarkable progress in developing complex reasoning abilities in Vision-Language Models (VLMs). Leading models, such as OpenAI GPT4o/GPT-o1 (Hurst et al., 2024; Jaech et al., 2024a), Gemini-2.5 (Team et al., 2023), VL-Rethinker (Wang et al., 2025) achieve superior performance on various multimodal reasoning benchmarks like MathVista (Lu et al., 2023), MMMU (Yue et al., 2024), MEGA-Bench (Chen et al., 2024), etc. A common paradigm underpinning these state-of-the-art VLMs involves processing multimodal queries to extract relevant cues, followed by a reasoning process (CoT (Wei et al., 2022)) conducted purely in the textual format.
@@ 2 Problem Formulation
$r(\mathbf{x},\mathbf{y})=\begin{cases}1&\text{if the solution }\mathbf{y}\text{ contains the correct answer to query }\mathbf{x},\\ 0&\text{otherwise}.\end{cases}$
@@ 3 Warm-Start Instruction Tuning
We aim to cultivate a novel pixel-space reasoning paradigm leveraging existing Visual-Language Models (VLMs). However, instruction-tuned models such as Qwen2.5-VL-Instruct exhibit limited zero-shot proficiency in executing novel visual operations (as shown in the analysis in the appendix), likely due to their absence in standard training data. To lay the groundwork for utilizing visual operations in subsequent reinforcement learning, we describe in this section our approach to data curation and instruction tuning.
@@ 4 Curiosity-Driven Reinforcement Learning
The warm-started model typically suffers from a disparity in its capabilities: proficient textual reasoning versus nascent pixel-space reasoning. This inherent imbalance creates a "learning trap" that impedes the development of pixel-space reasoning, stemming from two synergistic issues. Firstly, the model’s initial limited mastery over visual operations frequently leads to failure or incorrect outputs, resulting in a higher incidence of negative feedback compared to text-mediated reasoning. Secondly, a significant portion of training queries does not rigorously demand visual processing for a correct response, allowing the model to ignore the outcomes of visual operations or default to its stronger textual reasoning. This interplay fosters a detrimental cycle where initial failures discourage further attempts, leading to the premature abandonment of exploring and mastering visual operations. As shown in Fig. 6, when training the Warm-Start Model with standard RL  (DeepSeek-AI et al., 2025; Wang et al., 2025) without proper incentives, the policy learns to bypass the nascent visual operations.
@@
This reward scheme offers a dynamic reward mechanism that automatically tune the exploration bonus as training proceeds. As illustrated in Fig. 6, the curiosity bonus will naturally diminishes as the policy explores more on pixel-space reasoning. This prevents the policy policy from reward hacking – overly relying on the exploration bonus regardless of final correctness.
@@ 5 Experiments
[…]
@@ 5.1 Main Results
Table 1 shows that Pixel-Reasoner achieves the highest open-source results across all four benchmarks. Remarkably, Pixel-Reasoner, at a mere 7B parameters, not only surpasses substantially larger open-source models like the 27B Gemma3 across all benchmarks, but also outperforms specialized models that depend on external tools, such as IVM-Enhance (GPT-4V). Furthermore, Pixel-Reasoner’s exceptional capabilities extend to outperforming leading proprietary models, evidenced by its significant 5.1 percentage point lead over Gemini-2.5-Pro on V-star Bench (84.3 vs 79.2), and achieving the overall highest scores amongst all models listed. We observe that RL training is monumental to the great results. Our ablation model "Warm-Start Model (w/o RL)" has a lower performance than the original checkpoint Qwen2.5-VL on many benchmarks. After RL training, the performance skyrockets to SoTA level. This reflects the necessity of RL training to cultivate pixel-level reasoning.
@@
5.2 Key Factors for Cultivating Pixel-Space Reasoning
This section investigates two critical factors in fostering pixel-space reasoning through RL: first, the policy’s proficiency in utilizing visual operations lays the foundation for RL, and second, the role of incentives in encouraging the adoption of pixel-space reasoning during RL. To gain deeper insights into the results in Tab. 1, we analyze the training dynamics of various baselines, shown in Fig. 7. Specifically, the left panel shows the proportion of training rollouts that employ pixel-space reasoning strategies, while the middle panel indicates the error rate associated with the execution of visual operations. The right panel compares the expected correctness between the two reasoning paradigms, highlighting the disparity in the two capabilities over training time. In addition, the appendix provides concrete examples of trajectories that bypass pixel-space reasoning, illustrating "the learning trap."
@@ 5.2 Key Factors for Cultivating Pixel-Space Reasoning
  • The No-Correction Model (blue lines), trained solely on single-pass expert trajectories, initially exhibits an increase in RaPR, suggesting a propensity for attempting visual operations. However, this trend is quickly overshadowed by a significant and persistent rise in the failure rate of these operations. This elevated error rate points to the model’s inability to effectively respond to unexpected or erroneous outcomes from visual tasks. This deficiency stems directly from the absence of error-induced self-correction trajectories during its instruction-tuning phase. Consequently, the policy increasingly relies on pixel-space reasoning while simultaneously ignoring the outcomes of visual operations and favoring textual reasoning. Interestingly, we observe that the resulting reasoning trajectories involve error messages from visual operations but can still arrive at a correct answer. This indicates reward hacking: the policy earns curiosity bonus by superficially executing visual operations, and meanwhile it also earns correctness reward by essentially relying on textual reasoning to arrive at the final answer.
@@ 6 Related Work
Post-Training for Vision-Language Models. Post-training techniques, such as instruction tuning and reinforcement learning, are critical for adapting large Vision-Language Models (VLMs) to complex tasks beyond initial pre-training. LLaVA (Liu et al., 2023), Llava-OV (Li et al., 2024b), Infinity-MM (Gu et al., 2024), and MAmmoTH-VL (Guo et al., 2024) has shown that scaling instruction tuning datasets and increasing task diversity significantly enhances VLM generalization across various multimodal benchmarks.
Recently, a growing body of work applies RL to the multimodal domain (Deng et al., 2025; Huang et al., 2025; Feng et al., 2025). These approaches typically employ multi-stage pipelines, starting with SFT on costly data distillation and then applying RL to further refine the model’s reasoning capabilities. VL-Rethinker (Wang et al., 2025) investigates more direct RL approaches to foster slow-thinking in VLMs, and introduced selective sample replay (SSR) to counteract the vanishing advantages problem in GRPO.
Vision-Language Models with Tools. Recent research has explored augmenting VLMs with external tools or enabling them to perform pixel-level operations on inputs. Chain-of-Manipulation (Qi et al., 2025), Visual-Program-Distillation (VPD) (Hu et al., 2024b) focus on training models to effectively utilize tools or distill tool-based reasoning. Visual Sketchpad (Hu et al., 2024a) equips models, such as GPT-4o, with tools like depth perception and Python plotting. Models like o3 (Jaech et al., 2024b) demonstrate an ability to "think with images" by dynamically applying operations like zooming or flipping to improve visual understanding. Specific tools such as Instruction-Guided Visual Masking (Zheng et al., 2024) and visual guided search (Wu and Xie, 2024) has been integrated into these frameworks.
@@ 7 Conclusion
In this paper, we show how to incentivize the pixel-space reasoning from an existing vision-language models for the first time. Our warm-start instruction tuning and curiosity-driven RL are both essential to achieve the state-of-the-art performance. However, our work is currently still limited to two primary operations, which is insufficient for broader tasks. Our framework is easily extensible to other operations like depth map, image search, etc. In the future, the community can work together to enrich the visual operations to enhance the pixel-space reasoning in VLMs.
@@ Appendix B Derivations of Curiosity-Driven Reward
However, directly applying this standard formulation has two problems. Firstly, this formulation has an over-satisfaction issue. The term $-\lambda_{2}(\mathbf{n}_{\text{vo}}(\mathbf{y})-\mathrm{N})$ would provide a positive reward if $\mathbf{n}_{\text{vo}}(\mathbf{y})<\mathrm{N}$ (i.e., the constraint is "over-satisfied"), potentially encouraging the policy to use far fewer visual operations than necessary. Secondly, the term $-\lambda_{1}(\mathrm{H}-\text{RaPR}(\mathbf{x}))$ operates on the expectation-level and does not properly reward individual responses $y\sim\pi_{\theta}$.
@@ C.1 Protocols of Visual Operations
This operation allows the model to zoom in on a specific region of an image by providing a bounding box. The input includes a two-dimensional bounding box bbox_2d—a list of numeric coordinates $[x_{1},y_{1},x_{2},y_{2}]$ constrained within the image dimensions—and a target_image index indicating which image to operate on (indexed from 1, where 1 refers to the original image). This operation helps the model focus on fine-grained details.
@@ C.3 Training Details
Implementation Details. For Instruction Tuning, we adapt the Open-R1 code to implement SFT loss with loss masks. For RL, we implement based on OpenRLHF. We adopt GRPO [DeepSeek-AI et al., 2025] with selective sample replay [Wang et al., 2025], because we witness significant issues of vanishing advantages. As shown in Fig. 9, our reward scheme incorporates curiosity bonus and efficiency penalty in addition to correctness rewards, which provides more variance in rewards. However, the ratio of queries that suffer from reward uniformity steadily increases to 90% as training progresses, leading to a drastic plunge in performance evidenced by the ratios of "response-all-incorrect" queries. During RL training, we employed a near on-policy RL paradigm, where the behavior policy was synchronized with the improvement policy after every 512 queries, which we define as an episode. The replay buffer for SSR persisted for the duration of each episode before being cleared. For each query, we sampled 8 responses. The training batch size was set to 256 query-response pairs. Our 7B model is trained on $4\times 8$ sets of A800 (80G) for $20$ hours .
@@
Training Hyperparameters. For Instruction Tuning, we use a batch size of $128$. The learning rate is $1e^{-6}$ with 10% warm up steps. For RL, we set employ a cosine learning rate schedule with initial learning rate $1e^{-6}$ and 3% warm up iterations. During RL training, we sample 8 trajectories per training query and set hyperparameters to $\alpha=0.5$, $\beta=0.05$, $\mathrm{H}=0.3$, and $\mathrm{N}=1$. This configuration reflects our objectives: the threshold $\mathrm{H}=0.3$ encourages the policy to utilize pixel-space reasoning in approximately 30% of responses generated for a given query, while $\mathrm{N}=1$ promotes efficiency by favoring responses that require at most one visual operation. Under these parameters, a response can receive a maximum exploration bonus of approximately $0.5\times(0.3-1/8)\approx 0.0875$, while each additional visual operation beyond the first incurs a penalty of $-0.05$.

### v3

@@ Abstract
00footnotetext: Corresponding to: dlwlrma314516@gmail.com, jasper.whz@outlook.com, wenhuchen@uwaterloo.ca
@@ 1 Introduction
Recent advancements have demonstrated remarkable progress in developing complex reasoning abilities in Vision-Language Models (VLMs). Leading models achieve superior performance on various multimodal reasoning benchmarks like MathVista (Lu et al., 2023), MMMU (Yue et al., 2024), MEGA-Bench (Chen et al., 2024), etc. A common paradigm underpinning these state-of-the-art VLMs involves processing multimodal queries to extract relevant cues, followed by a reasoning process (CoT (Wei et al., 2022)) conducted purely in the textual format.
@@ 2 Problem Formulation
[…]
@@ 3 Warm-Start Instruction Tuning
We aim to cultivate a novel pixel-space reasoning paradigm leveraging existing Visual-Language Models (VLMs). However, a significant initial challenge is that instruction-tuned models such as Qwen2.5-VL-Instruct exhibit limited zero-shot proficiency in executing novel visual operations (as shown in the analysis in the appendix). This limitation likely stems from the absence of such interactive operations in their standard training data. To lay the necessary groundwork for utilizing visual operations in subsequent reinforcement learning, we describe in this section our approach to data curation and instruction tuning.
@@ 4 Curiosity-Driven Reinforcement Learning
The warm-started model typically suffers from a disparity in its capabilities: proficient textual reasoning versus nascent pixel-space reasoning. This inherent imbalance creates a "learning trap" that impedes the development of pixel-space reasoning, stemming from two synergistic issues. Firstly, the model’s initial limited mastery over visual operations frequently leads to failure or incorrect outputs, resulting in a higher incidence of negative feedback compared to text-mediated reasoning. Secondly, a significant portion of training queries does not rigorously demand visual processing for a correct response, allowing the model to ignore the outcomes of visual operations or default to its stronger textual reasoning. This interplay fosters a detrimental cycle where initial failures discourage further attempts, leading to the premature abandonment of exploring and mastering visual operations. As shown in Fig. 6, when training the Warm-Start Model with standard RL  (DeepSeek-AI et al., 2025; Wang et al., 2025a) without proper incentives, the policy learns to bypass the nascent visual operations.
@@
This reward scheme offers a dynamic reward mechanism that automatically tunes the exploration bonus as training proceeds. As illustrated in Fig. 6, the curiosity bonus is designed to be highest at the start of training when RaPR is low, actively incentivizing exploration. As the policy masters pixel-space reasoning, the bonus naturally diminishes. This dynamic property prevents the policy from "reward hacking"—a state where it might learn to overly rely on the exploration bonus for its reward, rather than optimizing for task correctness.
@@ 5 Experiments
Pixel-Reasoner was trained on $8\times\mathrm{A800(80G)}$ GPUs, using Open-R1 and OpenRLHF for instruction tuning and reinforcement learning respectively. We adopt GRPO (DeepSeek-AI et al., 2025) with selective sample relay due to vanishing advantages (Wang et al., 2025a). We include training details in the appendix, and will release code, models, and data to support reproducibility.
@@ 5.1 Main Results
Table 1 shows that Pixel-Reasoner achieves the highest open-source results across all four benchmarks. Remarkably, Pixel-Reasoner, at a mere 7B parameters, not only surpasses substantially larger open-source models like the 27B Gemma3 across all benchmarks, but also outperforms specialized models that depend on external tools, such as IVM-Enhance (GPT-4V). Furthermore, Pixel-Reasoner’s exceptional capabilities extend to outperforming leading proprietary models, evidenced by its significant 5.1 percentage point lead over Gemini-2.5-Pro on V-star Bench (84.3 vs 79.2), and achieving the overall highest scores amongst all models listed. We observe that RL training is monumental to the great results. Our ablation model "Warm-Start Model (w/o RL)" has a lower performance than the original checkpoint Qwen2.5-VL on many benchmarks. After RL training, the performance skyrockets to SoTA level. This reflects the necessity of RL training to cultivate pixel-level reasoning. The key advantage of the RL phase is its ability to overcome the "learning trap" identified in Section 4. Unlike supervised fine-tuning, which relies on static expert demonstrations, RL allows the model to actively explore the nascent pixel-space operations, and learn from trial-and-error. This exploration is critical for building a robust policy that does not default to its stronger, pre-existing textual reasoning pathways.
@@
5.2 Investigating the "Learning Trap": Key Factors for Cultivating Pixel-Space Reasoning
This section investigates the "learning trap" by analyzing two critical factors in fostering pixel-space reasoning through RL: first, the policy’s foundational proficiency in utilizing visual operations, and second, the role of incentives in encouraging the adoption of pixel-space reasoning during RL. To gain deeper insights into the results in Tab. 1, we analyze the training dynamics of various baselines, shown in Fig. 7. Specifically, the left panel shows the proportion of training rollouts that employ pixel-space reasoning strategies, while the middle panel indicates the error rate associated with the execution of visual operations. The right panel compares the expected correctness between the two reasoning paradigms, highlighting the disparity in the two capabilities over training time. In addition, the appendix provides concrete examples of trajectories that bypass pixel-space reasoning, illustrating "the learning trap."
@@ 5.2 Investigating the "Learning Trap": Key Factors for Cultivating Pixel-Space Reasoning
  • The No-Correction Model (blue lines), trained solely on single-pass expert trajectories, initially exhibits an increase in RaPR, suggesting a propensity for attempting visual operations. However, this trend is quickly overshadowed by a significant and persistent rise in the failure rate of these operations. This high error rate highlights the model’s inability to effectively respond to unexpected or erroneous outcomes from visual tasks. This deficiency stems directly from the absence of error-induced self-correction trajectories during its instruction-tuning phase. Consequently, the policy enters a reward-hacking loop: it earns the curiosity bonus by superficially executing visual operations (increasing its RaPR) but ignores the resulting errors, ultimately relying on its textual reasoning to arrive at the final answer and secure the correctness reward.
@@ 6 Related Work
Post-Training for Vision-Language Models. Post-training techniques, such as instruction tuning and reinforcement learning (Ouyang et al., 2022; Wang et al., 2025b, c), are critical for adapting large Vision-Language Models (VLMs) to complex tasks beyond initial pre-training. LLaVA (Liu et al., 2023), Llava-OV (Li et al., 2024b), Infinity-MM (Gu et al., 2024), and MAmmoTH-VL (Guo et al., 2024) has shown that scaling instruction tuning datasets and increasing task diversity significantly enhances VLM generalization across various multimodal benchmarks.
Recently, a growing body of work applies RL to the multimodal domain (Deng et al., 2025; Huang et al., 2025; Feng et al., 2025). These approaches typically employ multi-stage pipelines, starting with SFT on costly data distillation and then applying RL to further refine the model’s reasoning capabilities. VL-Rethinker (Wang et al., 2025a) investigates more direct RL approaches to foster slow-thinking in VLMs, and introduced selective sample replay (SSR) to counteract the vanishing advantages problem in GRPO.
Vision-Language Models with Tools. Recent research has explored augmenting VLMs with external tools or enabling them to perform pixel-level operations on inputs. Chain-of-Manipulation (Qi et al., 2025), Visual-Program-Distillation (VPD) (Hu et al., 2024b) focus on training models to effectively utilize tools or distill tool-based reasoning. Visual Sketchpad (Hu et al., 2024a) equips models, such as GPT-4o, with tools like depth perception and Python plotting. Models like o3 (Jaech et al., 2024) demonstrate an ability to "think with images" by dynamically applying operations like zooming or flipping to improve visual understanding. Specific tools such as Instruction-Guided Visual Masking (Zheng et al., 2024) and SEAL (Wu and Xie, 2024) has been integrated into these frameworks.
@@ 7 Conclusion
In this paper, we show how to incentivize the pixel-space reasoning from an existing vision-language models for the first time. Our warm-start instruction tuning and curiosity-driven RL are both essential to achieve the state-of-the-art performance. However, our work is currently still limited to two primary operations, which is insufficient for broader tasks. Our framework is easily extensible to other operations like depth map, image search, etc. In the future, the community can work together to not only enrich the library of visual operations but also to develop more sophisticated intrinsic rewarding mechanisms for pixel-space reasoning, enabling more powerful vision-language models.
@@ Appendix B Derivations of Curiosity-Driven Reward
However, directly applying this standard formulation has two problems. Firstly, this formulation has an over-satisfaction issue. The term $\lambda_{2}(\mathbf{n}_{\text{vo}}(\mathbf{y})-\mathrm{N})$ would provide a positive reward if $\mathbf{n}_{\text{vo}}(\mathbf{y})<\mathrm{N}$ (i.e., the constraint is "over-satisfied"), potentially encouraging the policy to use far fewer visual operations than necessary. Secondly, the term $\lambda_{1}(\mathrm{H}-\text{RaPR}(\mathbf{x}))$ operates on the expectation-level and does not properly reward individual responses $y\sim\pi_{\theta}$.
@@ C.1 Protocols of Visual Operations
This operation allows the model to zoom in on a specific region of an image by providing a bounding box. The input includes a two-dimensional bounding box bbox_2d—a list of numeric coordinates $x_{1},y_{1},x_{2},y_{2}]$ constrained within the image dimensions—and a target_image index indicating which image to operate on (indexed from 1, where 1 refers to the original image). This operation helps the model focus on fine-grained details.
@@ C.3 Training Details
Implementation Details. For Instruction Tuning, we adapt the Open-R1 code to implement SFT loss with loss masks. For RL, we implement based on OpenRLHF. We adopt GRPO [DeepSeek-AI et al., 2025] with selective sample replay [Wang et al., 2025a], because we witness significant issues of vanishing advantages. As shown in Fig. 9, our reward scheme incorporates curiosity bonus and efficiency penalty in addition to correctness rewards, which provides more variance in rewards. However, the ratio of queries that suffer from reward uniformity steadily increases to 90% as training progresses, leading to a drastic plunge in performance evidenced by the ratios of "response-all-incorrect" queries. During RL training, we employed a near on-policy RL paradigm, where the behavior policy was synchronized with the improvement policy after every 512 queries, which we define as an episode. The replay buffer for SSR persisted for the duration of each episode before being cleared. For each query, we sampled 8 responses. The training batch size was set to 256 query-response pairs. Our 7B model is trained on $4\times 8$ sets of A800 (80G) for $20$ hours .
@@
Training Hyperparameters. For Instruction Tuning, we use a batch size of $128$. The learning rate is $1e^{-6}$ with 10% warm up steps. For RL, we set employ a cosine learning rate schedule with initial learning rate $1e^{-6}$ and 3% warm up iterations. During RL training, we sample 8 trajectories per training query and set hyperparameters to $\alpha=0.5$, $\beta=0.05$, $\mathrm{H}=0.3$, and $\mathrm{N}=1$. This configuration reflects our objectives: the threshold $\mathrm{H}=0.3$ encourages the policy to utilize pixel-space reasoning in approximately 30% of responses generated for a given query, while $\mathrm{N}=1$ promotes efficiency by favoring responses that require at most one visual operation. Under these parameters, a response can receive a maximum exploration bonus of approximately $0.5\times(0.3-1/8)\approx 0.0875$, while each additional visual operation beyond the first incurs a penalty of 0.05-0.05.